OpenStudy (anonymous):

The sum of the digits of a two-digit number is 12. The number formed by interchanging the digits is 54 more than the original number. What is the original number? 39 48 57

OpenStudy (mertsj):

TU = original number = 10T+U UT = number with digits interchanged = 10U+T 10U+T=10T+U+54 T+U=12 Solve the system

OpenStudy (anonymous):

48

OpenStudy (mertsj):

If you add 54 to 48 do you get 84?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

none of the answers would add with54 to get 84

OpenStudy (mertsj):